# On Speeding Up Language Model Evaluation

### v2

@@ 3.2 Algorithm 2 – UCB-E with Low-Rank Factorization (UCB-E-LRF)
Algorithm 1, the UCB-E algorithm, treats each method independently and samples examples for a method uniformly at random. In practice, however, methods and examples are correlated with each other and we can predict the outcome of a particular method-example pair from previous observations. Being able to accurately predict the outcome gives better estimate of the score of a method. Additionally, instead of sampling examples uniformly at random, prioritizing examples with large prediction uncertainties reduces the overall uncertainty in the scores of each method, which provides more accurate score estimates in the subsequent steps. In this section, we propose a low-rank factorization based algorithm (UCB-E-LRF) to capture the above intuition.
@@
A natural approach to modeling these method-example interactions is through low-rank matrix factorization. Intuitively, if the method-examples are very correlated, there should exist $U\in\mathbb{R}^{n\times r}$ and $V\in\mathbb{R}^{m\times r}$ and rank $r\ll\min(n,m)$ such that $E\approx UV^{\top}$. Obtaining $U$ and $V$ from observed scoring matrix $E^{\rm obs}$ allows us to estimate the scores of all method-example pairs from just a few scores. For instance, if the matrix $E$ were an exact rank-1 matrix, only $n+m-1$ scores would be needed to recover the full scoring matrix $E$, which is far more efficient than exhaustively evaluating all $n\times m$ combinations. To find $U$ and $V$, we use alternating least squares [26] to optimize the following problem:
$\min_{U,V}\Big{|}\Big{|}O\odot\Big{(}UV^{T}-E^{\rm obs}\Big{)}\Big{|}\Big{|}_{F}^{2}.$ (1)
@@
We can then combine the observed matrix $E^{\rm obs}$ and estimated matrix $\hat{E}=UV^{T}$, using $O$ as a gating function, to arrive at a prediction of the score $\mu_{i}$ as follows:
$\hat{\mu}_{i}=\frac{1}{m}\sum_{j=1}^{m}O_{i,j}E^{\rm obs}_{i,j}+(1-O_{i,j})\hat{E}_{i,j}.$ (2)
@@
The score estimation gives us a better way to estimate score of every method by exploiting the correlation between methods and examples at every step. We now discuss how we estimate prediction uncertainties with ensembles.
By ensembling the factorization model, we obtain an empirical distribution of predictions for each method-example pair. Consequently, the standard deviation of the predictions can be used to quantify uncertainty. If the predictions are very similar to each other, this suggests the ensembles tend to agree with each other, and hence there is little uncertainty. In order to achieve sufficient model variance to approximate uncertainty, we uniformly at random drop some values from $E^{\rm obs}$ when solving the factorization problem in Equation 1 for each model  [23, 67]. For an ensemble of size $S$ the estimate for $\hat{E}$ becomes the following:
$\hat{E}=\frac{1}{S}\sum_{s=1}^{S}U^{(s)}(V^{(s)})^{T}.$ (3)
Then the uncertainty matrix $R\in\mathbb{R}^{n\times m}$ is given by:
$R=\sqrt{\frac{1}{S}\sum_{s=1}^{S}\Bigg{[}\Big{(}1-O\Big{)}\odot\Big{(}U^{(s)}(V^{(s)})^{T}-\hat{E}\Big{)}\Bigg{]}^{2}}.$ (4)
Adapative selection.
Bringing score estimation and uncertainty estimation together, we propose our second algorithm UCB-E-LRF ($\mathcal{A}_{\rm uel}$; Algorithm 2), an extension of Algorithm 1 leveraging the low-rank factorization model $\mathcal{M}$. The high level structure of the algorithms are very similar to each other. Specifically, we estimate the upper confidence bound (UCB) for each method based on a linear combination of score and uncertainty estimation, select the method with the greatest UCB, and evaluate the method on an example that has the highest uncertainty. Lastly, we introduce a warm-up budget $T_{0}$. This is because $\mathcal{M}$ requires a minimum number of initial observations in order to estimate the low-rank representation accurately. Therefore, before the active selection phase (line 3-9 in Algorithm 2), we have a warm-up phase (line 1-2 in Algorithm 2), where $T_{0}$ method-example pairs are sample uniformly at random to evaluate before moving on to the “active" phase.
Algorithm 2 UCB-E-LRF ($\mathcal{A}_{\rm uel}(T,\mathcal{F},\mathcal{X};\mathcal{M},r,S,T_{0},\eta)$)
Input: The evaluation budget $T$, a set of methods $\mathcal{F}$, a set of examples $\mathcal{X}$, the low-rank factorization model $\mathcal{M}$, the rank $r$, the ensemble size $S$, the warm-up budget $T_{0}$, the uncertainty scaling $\eta$.
@@
1:  Uniformly draw $T_{0}$ method-example pairs from $[n]\times[m]$ and get the observation matrix $O\in\{0,1\}^{n\times m}$ and observed scoring matrix $E^{\rm obs}\in\left([0,1]\cup\{?\}\right)^{n\times m}$ w.r.t. these $T_{0}$ evaluations.
2:  $\hat{E},R\leftarrow\mathcal{M}(E^{\rm obs},O;r,S)$; $\forall f_{i}\in\mathcal{F},B_{i}\leftarrow\frac{1}{m}\sum_{j=1}^{m}(O_{ij}E^{\rm obs}_{ij}+(1-O_{ij})\hat{E}_{ij}+\eta R_{ij})$
@@
4:     Select: Draw uniformly at random $i\in\arg\max_{k\,\mid\,(\sum_{j=1}^{m}O_{kj})\neq m}B_{k}$; Draw uniformly at random $j\in\arg\max_{k\,\mid\,O_{ik}=0}R_{ik}$.
5:     Evaluate: Run inference for the method-example pair $(f_{i},x_{j})$, score the result, and receive $e(f_{i}(x_{j}))$; $E^{\rm obs}_{ij}\leftarrow e(f_{i}(x_{j}))$.
6:     Update: $O_{i,j}\leftarrow 1$; $\hat{E},R\leftarrow\mathcal{M}(E^{\rm obs},O;r,S)$; $\forall f_{i}\in\mathcal{F},B_{i}\leftarrow\frac{1}{m}\sum_{j=1}^{m}(O_{ij}E^{\rm obs}_{ij}+(1-O_{ij})\hat{E}_{ij}+\eta R_{ij})$.
@@ 4 Experiments
In this section, we study the performance of UCB-E and UCB-E-LRF in a variety of real settings. We begin by describing our dataset and metrics, and then analyze the experimental results and ablations.
Figure 3: Comparison of baselines and our proposed algorithms on six datasets evaluated with various top 1 precision and NDCG@10. The vertical axes of all plots represent performance of a metric and horizontal axies of all plots represent the percentage of method-example pairs evaluated. All results are aggregated based on 50 trials with different random seeds. Our proposed algorithms: UCB-E and UCB-E-LRF consistently require much less evaluation to achieve the same performance as baselines.
@@ 4.1 Datasets
To assess the performance of our algorithms under a variety of use cases, we test with three datasets AlpacaEval [44], Grade School Math 8K (GSM8K) [18] and Physical Interaction: Question Answering (PIQA) [3], together with different settings of the method set $\mathcal{F}$ and the scoring function $e$; Table 2 summarizes the statistics of each dataset along with what method $\mathcal{F}$ and scoring function $e$. For AlpacaEval, we design two sets of $\mathcal{F}$. The first set $\mathcal{F}$ contains all LLMs reported by [44]. The second $\mathcal{F}$ contains the same LLMs except for the annotator model GPT-4 Turbo. The latter $\mathcal{F}$ that does not contain GPT-4 Turbo makes the learning more challenging and interesting since the annotator, GPT-4 Turbo, is a clear winner among all LLMs. For GSM8K and PIQA, we set $\mathcal{F}$ as a set of prompts to simulate prompt engineering, or a set of LLMs with various sampling configurations, to mimic model selection and hyperparameter tuning. We provide more information about these datasets in Appendix A.
Table 2: Statistics of each dataset used in our experiments. The examples $\mathcal{X}$ of each dataset are the questions from their respective evaluation benchmark. More details of the datasets can be found in Appendix A.
Dataset Name Size $n\times m$ Method $\mathcal{F}$ Scoring Function $e$ $H_{1}$
AlpacaEval $154\times 805$ Various LLMs GPT4-turbo annotator 966
AlpacaEval (Drop Annotator) $153\times 805$ Various LLMs excluding GPT4-turbo GPT4-turbo annotator 4462
GSM8K Prompts $205\times 784$ Mistral-7B [32] with different prompts regex match with correct answer 107445
GSM8K Models $122\times 1000$ Various LLMs and sampling configurations regex match with correct answer 20562
PIQA Prompts $177\times 1546$ Tulu-7B [59] with different prompts regex match with correct choice 66284
PIQA Models $103\times 1000$ Various LLMs and sampling configurations regex match with correct choice 10273
@@ 4.2 Metrics
To determine if an algorithm finds the best method, we can directly check if the algorithm’s prediction matches with the best method from our empirical data. However, because every method is empirically evaluated on limited number of examples, it is possible that one method has slightly lower average performance than another method whereas if we were to evaluate on more examples, the former would achieve a higher performance. To this end, we calculate top 1 precision by first determining a set of methods we consider equally good and we check if the predicted method from an algorithm is a member of that set. We propose two ways to determine if a set of methods are equally good: performance gap and statistical significance. For performance gap $\epsilon$ top 1 precision, we consider all methods whose performance is within $\epsilon$ of the best empirical method to be equally good. For the statistical significance $p$ top 1 precision, we perform McNemar’s statistical test [47] for each method against the best empirical method. If we cannot reject the null hypothesis that the performance of one method is the same as the best empirical method up to a significance level $p$, then that method is considered equally good as the best method. We evaluate all baselines and our algorithms on two $\epsilon$ values $\{0.001,0.01\}$, and two $p$ values $\{0.01,0.1\}$ simulating a diverse need of precision.
@@
Although our focus is to identify the best method quickly, it is sometimes also desirable to generally rank top K promising methods high. We therefore evaluate with normalized discounted cumulative gain (NDCG) at K. NDCG@K takes as input the top K prediction by an algorithm and the higher their true ranks are, the higher the NDCG is. We choose $K=10$ to evaluate all algorithms.
@@ 4.3 Set-Up of Our algorithms and Baselines
We introduce the set-up of our two algorithms, UCB-E and UCB-LRF, together with three baselines, Row Mean Imputation, Filled Subset and LRF.
UCB-E.
Our proposed algorithm as shown in Algorithm 1. We use $a=1$ since this consistently yields the best performance across all datasets. This parameter is ablated in Figure 7 of Appendix B.
UCB-E-LRF.
Our proposed algorithm as shown in Algorithm 2. For all datasets, we use rank $r=1$ for low rank factorization with an ensemble size of $S=64$. We use $5\%$ of data for warm up i.e. $T_{0}=0.05\times n\times m$ and $\eta=5$.
Row Mean Imputation.
In this baseline, we select method-example pairs uniformly at random from all pairs. The score for each method is calculated as the average of all the scores evaluated so far for that method. The detailed algorithmic description is in Algorithm 3.
Filled Subset.
Instead of randomly selecting method-example pairs, filled subset first selects an example index $j$ uniformly at random. Then all methods are evaluated on example $j$. If all methods have been evaluated, the algorithm selects a new example index from remaining ones. The score for each method is calculated as the average of all the scores evaluated so far for that method. The detailed algorithmic description is in Algorithm 4. Although row mean imputation and filled subset do not have learning component, they both produce an unbiased estimate of the real score for each method at all time.
LRF.
Similar to row mean imputation, LRF (low rank factorization) randomly selects method-example pairs to evaluate. For estimating the score of each method, LRF calculates the average of all scores for that method. That is, for evaluated examples, the score is the actual observed score, and for examples yet to be evaluated, the score is the estimated score from LRF. The detailed algorithmic description is in Algorithm 5.
Finally, for better practicality and efficiency, all algorithms and baselines are implemented with a batch size $b$. That is, instead of selecting and evaluating just one example at a time, the algorithms select $b$ examples. For algorithms that involve randomly drawing an example, they are modified to randomly drawing $b$ examples without replacement. For algorithms that draw an example with the $\arg\max$ operator, they are modified to drawing with the $\mathrm{Top-}b$ operator. We use $b=32$ for reporting all main results.

### v3

@@ 3.2 Algorithm 2 – UCB-E with Low-Rank Factorization (UCB-E-LRF)
Algorithm 1, the UCB-E algorithm, treats each method independently and samples examples for a method uniformly at random. In practice, however, methods and examples are correlated with each other and we can predict the outcome of a particular method-example pair from previous observations. Being able to accurately predict the outcome gives a better estimate of the score of a method. Additionally, instead of sampling examples uniformly at random, prioritizing examples with large prediction uncertainties reduces the overall uncertainty in the scores of each method, which provides more accurate score estimates in the subsequent steps. In this section, we propose a low-rank factorization based algorithm (UCB-E-LRF) to capture the above intuition.
@@
A natural approach to modeling these method-example interactions is through low-rank matrix factorization. Intuitively, if the method-examples are very correlated, there should exist $U\in\mathbb{R}^{m\times r}$ and $V\in\mathbb{R}^{n\times r}$ and rank $r\ll\min(m,n)$ such that $S\approx UV^{\top}$ (Candes and Recht, 2012; Chen et al., 2020; Cai et al., 2019). Obtaining $U$ and $V$ from the observed scoring matrix $S^{\rm obs}$ allows us to estimate the scores of all method-example pairs from just a few scores. For instance, if the matrix $S$ were an exact rank-1 matrix, only $m+n-1$ scores would be needed to recover the full scoring matrix $S$, which is far more efficient than exhaustively evaluating all $m\times n$ combinations. To find $U$ and $V$, we use alternating least squares (Hastie et al., 2015) to optimize the following problem:
$\min_{U,V}\Big{|}\Big{|}O\odot\Big{(}UV^{T}-S^{\rm obs}\Big{)}\Big{|}\Big{|}_{F}^{2}.$ (1)
@@
We can then combine the observed matrix $S^{\rm obs}$ and estimated matrix $\hat{S}=UV^{T}$, using $O$ as a gating function, to arrive at a prediction of the score $\mu_{i}$ as follows:
$\hat{\mu}_{i}=\frac{1}{n}\sum_{j=1}^{n}O_{ij}S^{\rm obs}_{ij}+(1-O_{ij})\hat{S}_{ij}.$ (2)
@@
The score estimation gives us a better way to estimate the score of every method by exploiting the correlation between methods and examples at every step. We now discuss how we estimate prediction uncertainties with ensembles.
By fitting many factorization models, we obtain an empirical distribution of predictions for each method-example pair. Consequently, the standard deviation of the predictions can be used to quantify uncertainty. If the predictions are very similar to each other, this suggests the ensembles tend to agree with each other, and hence there is little uncertainty. In order to achieve sufficient model variance to approximate uncertainty, we uniformly at random drop some values from $S^{\rm obs}$ when solving the factorization problem in Equation 1 for each model  (Efron and Tibshirani, 1994; Zoubir and Boashash, 1998). For an ensemble of size $C$ the estimate for $\hat{S}$ becomes the following:
$\hat{S}=\frac{1}{C}\sum_{c=1}^{C}U^{(c)}(V^{(c)})^{T}.$ (3)
Then the uncertainty matrix $R\in\mathbb{R}^{m\times n}$ is given by:
$R=\sqrt{\frac{1}{C}\sum_{c=1}^{C}\Bigg{[}\Big{(}1-O\Big{)}\odot\Big{(}U^{(c)}(V^{(c)})^{T}-\hat{S}\Big{)}\Bigg{]}^{2}}.$ (4)
Adaptive selection.
Bringing score estimation and uncertainty estimation together, we propose our second algorithm UCB-E-LRF ($\mathcal{A}_{\rm uel}$; Algorithm 2), an extension of Algorithm 1 leveraging the low-rank factorization model $\mathcal{M}$. The high level structure of the algorithms are very similar to each other. Specifically, we estimate the upper confidence bound (UCB) for each method based on a linear combination of score and uncertainty estimation, select the method with the greatest UCB, and evaluate the method on an example that has the highest uncertainty. Lastly, we introduce a warm-up budget $T_{0}$. This is because $\mathcal{M}$ requires a minimum number of initial observations in order to estimate the low-rank representation accurately. Therefore, before the active selection phase (line 3-9 in Algorithm 2), we have a warm-up phase (line 1-2 in Algorithm 2), where $T_{0}$ method-example pairs are sample uniformly at random to evaluate before moving on to the “active” phase.
Algorithm 2 UCB-E-LRF ($\mathcal{A}_{\rm uel}(T,\mathcal{F},\mathcal{X};\mathcal{M},r,C,T_{0},\eta)$)
Input: The evaluation budget $T$, a set of methods $\mathcal{F}$, a set of examples $\mathcal{X}$, the low-rank factorization model $\mathcal{M}$, the rank $r$, the ensemble size $C$, the warm-up budget $T_{0}$, the uncertainty scaling $\eta$.
@@
1:  Uniformly draw $T_{0}$ method-example pairs from $[m]\times[n]$ and get the observation matrix $O\in\{0,1\}^{m\times n}$ and observed scoring matrix $S^{\rm obs}\in\left([0,1]\cup\{?\}\right)^{m\times n}$ w.r.t. these $T_{0}$ evaluations.
2:  $\hat{S},R\leftarrow\mathcal{M}(S^{\rm obs},O;r,C)$; $\forall f_{i}\in\mathcal{F},B_{i}\leftarrow\frac{1}{n}\sum_{j=1}^{n}(O_{ij}S^{\rm obs}_{ij}+(1-O_{ij})\hat{S}_{ij}+\eta R_{ij})$
@@
4:     Select: Draw $i\in\arg\max_{k\,\mid\,(\sum_{j=1}^{n}O_{kj})\neq n}B_{k}$; Draw $j\in\arg\max_{k\,\mid\,O_{ik}=0}R_{ik}$.
5:     Evaluate: Run inference for the method-example pair $(f_{i},x_{j})$, score the result, and receive $s(f_{i}(x_{j}))$; $S^{\rm obs}_{ij}\leftarrow s(f_{i}(x_{j}))$.
6:     Update: $O_{ij}\leftarrow 1$; $\hat{S},R\leftarrow\mathcal{M}(S^{\rm obs},O;r,C)$; $\forall f_{i}\in\mathcal{F},B_{i}\leftarrow\frac{1}{n}\sum_{j=1}^{n}(O_{ij}S^{\rm obs}_{ij}+(1-O_{ij})\hat{S}_{ij}+\eta R_{ij})$.
@@ 4 Experiments
Figure 3: Comparison of algorithms on six datasets evaluated with various metrics. The vertical axes represents performance of a metric and horizontal axes represents the percentage of method-example pairs evaluated. All results are aggregated based on 50 trials with different random seeds. The datasets are ordered by decreasing $H_{1}$ (see Section 3.1). Larger $H_{1}$ values indicate settings where finding the best method is more difficult. Our proposed algorithms: UCB-E and UCB-E-LRF consistently require much less evaluation to achieve the same performance as baselines.
@@ 4.1 Datasets
Table 2: Statistics of each dataset used in our experiments. The examples $\mathcal{X}$ in each dataset correspond to the questions from their respective evaluation benchmarks. $H_{1}$ as defined in Section 3.1, reflects the difficulty of identifying the best method, with larger values indicating more challenging settings. More dataset details can be found in Appendix C. DA stands for Drop Annotator.
Method Set $\mathcal{F}$ Dataset Name Size $m\times n$ Scoring Function $s$ $H_{1}$
Various LLMs AlpacaEval $154\times 805$ GPT4-turbo annotator 966
Various LLMs excluding GPT4-turbo AlpacaEval (DA) $153\times 805$ GPT4-turbo annotator 4462
Mistral-7B with different prompts GSM8K Prompts $205\times 784$ regex match with correct answer 107445
Various LLMs and sampling configurations GSM8K Models $122\times 1000$ regex match with correct answer 20562
Tulu-7B with different prompts PIQA Prompts $177\times 1546$ regex match with correct choice 66284
Various LLMs and sampling configurations PIQA Models $103\times 1000$ regex match with correct choice 10273
To assess the performance of our algorithms under a variety of use cases, we test with three datasets AlpacaEval (Li et al., 2023), Grade School Math 8K (GSM8K) (Cobbe et al., 2021) and Physical Interaction: Question Answering (PIQA) (Bisk et al., 2020), together with different settings of the method set $\mathcal{F}$ and the scoring function $s$; Table 2 summarizes each dataset. For AlpacaEval, we design two sets of $\mathcal{F}$. The first set $\mathcal{F}$ contains all LLMs reported by Li et al. (2023). The second $\mathcal{F}$ contains the same LLMs except for the annotator model GPT-4 Turbo. The latter $\mathcal{F}$ that does not contain GPT-4 Turbo makes the learning more challenging and interesting since the annotator, GPT-4 Turbo, is a clear winner among all LLMs. For GSM8K and PIQA, we set $\mathcal{F}$ as a set of prompts to simulate prompt engineering, or a set of LLMs with various sampling configurations, to mimic model selection and hyperparameter tuning. We also observe that the datasets indeed exhibit low-rank properties as seen from ratios of singular values (Table 3) and explained variance of each principal component (Figure 6) for the data matrices. More information about these datasets is in Appendix C.
@@ 4.2 Metrics
To determine if an algorithm finds the best method, we can directly check if the algorithm’s prediction matches with the best method from our empirical data. However, because every method is empirically evaluated on a limited number of examples, it is possible that one method has a slightly lower average performance than another method whereas if we were to evaluate on more examples, the former would achieve a higher performance. To this end, we calculate top 1 precision by first determining a set of methods we consider equally good and checking if the predicted best method from an algorithm is a member of that set. We propose two ways to determine if a set of methods are equally good: performance gap and statistical significance. For performance gap $\epsilon$ top 1 precision, we consider all methods whose performance is within $\epsilon$ of the best empirical method to be equally good. For the statistical significance $p$ top 1 precision, we perform McNemar’s statistical test (McNemar, 1947) for each method against the best empirical method. If we cannot reject the null hypothesis that the performance of one method is the same as the best empirical method up to a significance level $p$, then that method is considered equally good as the best method. We evaluate all baselines and our algorithms on two $\epsilon$ values $\{0.001,0.01\}$, and two $p$ values $\{0.01,0.1\}$ simulating a diverse need of precision.
@@
Although our focus is to identify the best method quickly, it is sometimes also desirable to generally rank the top K promising methods high. We therefore evaluate with normalized discounted cumulative gain (NDCG) at K. NDCG@K takes as input the top K prediction by an algorithm and the higher their true ranks are, the higher the NDCG is. We choose $K=10$ to evaluate all algorithms.
@@ 4.3 Set-Up of Our algorithms and Baselines
UCB-E: We use $a=1$ since this consistently yields the best performance across all datasets. This parameter is ablated in Figure 7 of Appendix D. UCB-E-LRF: For all datasets, we use rank $r=1$ for low rank factorization with an ensemble size of $C=64$. We use $5\%$ of data for warm up i.e. $T_{0}=0.05\times m\times n$ and $\eta=5$. Row Mean Imputation: We select method-example pairs uniformly at random from all pairs. The score for each method is calculated as the average of all the scores evaluated so far for that method. Filled Subset: Instead of randomly selecting method-example pairs, filled subset first selects an example index $j$ uniformly at random. Then all methods are evaluated on example $j$. If all methods have been evaluated, the algorithm selects a new example index from remaining ones. The score for each method is calculated as the average of all the scores evaluated so far for that method. Although row mean imputation and filled subset do not have learning component, they both produce an unbiased estimate of the real score for each method at all times. LRF: Similar to row mean imputation, LRF (low rank factorization) randomly selects method-example pairs to evaluate. For estimating the score of each method, LRF calculates the average of all scores for that method. That is, for evaluated examples, the score is the actual observed score, and for examples yet to be evaluated, the score is the estimated score from LRF. The details for three baselines Row Mean Imputation, Filled Subset, and LRF are found in Appendix D Algorithm 3, 4, and 5, respectively.
Finally, to take advantaged of the parallelism available in modern computing hardware, all algorithms and baselines are implemented with a batch size $b$. For algorithms that randomly draw an example, they instead randomly draw $b$ examples without replacement. For algorithms that draw an example with the $\arg\max$ operator, they instead draw with the $\mathrm{Top-}b$ operator. We use $b=32$ for all experiments unless explicitly stated otherwise.